OpenStudy (hpfan101):

Find the point on the curve \[y^3 -4xy+x^2 =1\] inside Quadrant I having x=1.

OpenStudy (anonymous):

What happens of you make x=1 and rewrite the equation

OpenStudy (anonymous):

you there

OpenStudy (hpfan101):

\[y^3-4(1)y+(1)^2=1 \rightarrow y^3-4y+1=1 \rightarrow y^3-4y=0\rightarrow y(y^2-4)=0\]

OpenStudy (hpfan101):

So y would equal 0, 2, and -2

OpenStudy (anonymous):

but which one would be in Q-1

OpenStudy (hpfan101):

y=2 for sure. But would y=0 also be part of quadrant 1?

OpenStudy (anonymous):

I don't think so. Good work

OpenStudy (hpfan101):

Alright, thank you for the help!